OpenStudy (anonymous):

Graphing functions. -4x + y = -12 Please explain how you got the answer and show your steps using the equation button. Thank you!

OpenStudy (precal):

solve for y then It will be easy to show how to graph using the y intercept and the slope method

OpenStudy (anonymous):

@precal, But I don't know how to solve for y, that's why I submitted the question. >.<

OpenStudy (johnweldon1993):

You would ISOLATE 'y' ...so get it by itself on that left side -4x + y = -12 add 4x to both sides...this way it will cancel out on that left side -4x + y = -12 +4x +4x y = 4x - 12 This is a good starting point From here (slope intercept form) you have the slope, and the y-intercept y = mx + b 'm' is the slope 'b' is the y-intercept So when you compare that...to your equation you have slope = 4 y- intercept = -12 can you graph this from here?

OpenStudy (precal):

plot (0,-12) from that point you go up 4 and right 1 and plot another point go back to (0,-12) then you go down 4 and left 1 and plot another point that plots at least 3 points for you.